4. An electron in a cathode ray tube (CRT) accelerates from 2x104

 m/s to 6x106

 m/s over 1.50 cm.

(a) How long does the electron take to travel this 1.5 cm?

(b) What is its acceleration


1
Expert's answer
2021-07-21T09:06:09-0400

a)


"\frac{v_f+v_i}{2}t=s\\\frac{2\cdot10^4+6\cdot10^6}{2}t=0.015\\t=5.0\cdot10^{-9}\ s"

b)


"a=\frac{-2\cdot10^4+6\cdot10^6}{5.0\cdot10^{-9}}=1.2\cdot10^{15}\frac{m}{s^2}"